The lenght of a rectangle is 3 inches greater than the with. write a polynomial that represents the are of the rectangle
Answer by algebrahouse.com(1659)

"The length of a rectangle is 3 inches greater than the width. write a polynomial that represents the are of the rectangle"

x = width
x + 3 = length {length is 3 greater than width}

Area of a rectangle = width x length
= x(x + 3) {width x length}
= x^2 + 3x {used distributive property}
www.algebrahouse.com